OpenStudy (anonymous):

Water is lifted out of a well 20.0m deep by a motor rated at 1.60hp. Assuming efficiency 90{\rm \\%} , how many kilograms of water can be lifted in 1 min?

OpenStudy (shamim):

Required formula is p=w/t

OpenStudy (shamim):

Here p=effective power

OpenStudy (shamim):

W=work t=time

OpenStudy (shamim):

Given p=90%*1.6hp=(90/100)*1.6*746watt W=mgh=m*9.8*20 t=60sec m=?

OpenStudy (shamim):

Response plz!!!!!!!!

OpenStudy (anonymous):

thats math.... nextime post it in the math section ;)

OpenStudy (shamim):

That is math of physics??!!!